OpenStudy (anonymous):

Given the parent functions f(x) = log2 (3x − 9) and g(x) = log2 (x − 3), what is f(x) − g(x)? f(x) − g(x) = log2 (2x − 6) f(x) − g(x) = log2 (2x − 12) f(x) − g(x) = log2 one third f(x) − g(x) = log2 3

OpenStudy (♪chibiterasu):

Hmm, I'm not so sure about this one. I don't remember much about these types of functions from Algebra 1 Dx

OpenStudy (anonymous):

this is Alg2

OpenStudy (♪chibiterasu):

Im taking Geometry, so I am not at this level quite yet :c

OpenStudy (anonymous):

lol its ok

OpenStudy (anonymous):

f(x) − g(x) = log2 (2x − 6) f(x) − g(x) = log2 (2x − 12) f(x) − g(x) = log2 one third f(x) − g(x) = log2 3 So my answer is D

OpenStudy (anonymous):

f(x) = log2 (3x − 9) and g(x) = log2 (x − 3), what is f(x) − g(x) f(x) − g(x) = log2 (3x − 9) -log2 (x − 3) \[f(x)−g(x)=\log_2\frac{(3x−9)}{(x−3)}=\log_2 \frac{3(x−3)}{(x−3)} =\log_2 3\] Thus \[f(x) − g(x) = \log_2 3\] is the required answer

OpenStudy (anonymous):

@mgarrick77

OpenStudy (anonymous):

thank you soooooo much !

OpenStudy (anonymous):

@mgarrick77 medal please